OpenStudy (anonymous):

Simplify: 6a – 5b – 7b + 9a – 8a + 2

OpenStudy (anonymous):

once again combine like terms. you have \[6a+9a-8a=7a\] \[-5b-7b=-12b\] and so get \[7a-12b-2\]

OpenStudy (anonymous):

you should really do this in your head. it amounts to adding and subtracting

OpenStudy (anonymous):

as in \[5+9-8=7\] \[-5-7=-12\]

OpenStudy (anonymous):

imm in summer school dont have time aha

OpenStudy (anonymous):

i meant \[6+9-7\] summer school ick

OpenStudy (anonymous):

mind helping me a bit

OpenStudy (anonymous):

not at all

OpenStudy (anonymous):

Evaluate the following expression for r = 5 and q = -6 5r – (3r + 2q) + 4(2 – q)

OpenStudy (vishweshshrimali5):

satellite u are right

OpenStudy (anonymous):

satellite is amazing

OpenStudy (anonymous):

your job is to compute \[5\times 5-(3\times 5+2\times (-6))+4(2-(-6))\]

OpenStudy (anonymous):

someone must really hate you to make you do this. i would use a calculator

OpenStudy (anonymous):

they dont allow us to use them in ss idk why

OpenStudy (anonymous):

or not. just do \[25-(15-12)+4\times 8\] \[25-3+32\] \[23+32\] \[55\] check my arithmetic

OpenStudy (anonymous):

because, as i said, they hate you.

OpenStudy (anonymous):

so the answer is... lol

OpenStudy (anonymous):

i got 55 you want me to check it again?

OpenStudy (anonymous):

Evaluate the following expression for r = 5 and q = -6 5r – (3r + 2q) + 4(2 – q) shouldnt there be 2 answers??

OpenStudy (anonymous):

or just one?